Using LaGrange's method, solve the differential equations

p - qyIn(y) = zIn(y)


1
Expert's answer
2020-10-16T15:31:21-0400

"\displaystyle\textsf{The given partial differential equation is}\\
 p - q\, y\ln(y) = z\ln(y)\\


\mathrm{d}z =\frac{\partial z}{\partial x} \mathrm{d}x + \frac{\partial z}{\partial y} \mathrm{d}y \\

\mathrm{d}x = 1, \mathrm{d}y = -y\ln(y), \mathrm{d}z =z\ln(y) \\

\textsf{The Lagrange’s auxiliary equations}\\\textsf{for the given PDE is}\\

\frac{\mathrm{d}x}{1} = \frac{\mathrm{d}y}{-y\ln(y)} = \frac{\mathrm{d}z}{z\ln(y)}\\

\textsf{Choosing}\, \left(0, \frac{1}{y}, \frac{1}{z}\right)\, \textsf{as multipliers, we have}\\

\frac{1}{y}\mathrm{d}y + \frac{1}{z}\mathrm{d}z = 0 \\

\textsf{Integrating both sides, we have} \\

\int \frac{1}{y}\,\mathrm{d}y - \int \frac{1}{z}\,\mathrm{d}z= C_1 \\

\ln(y) - \ln(z) = C_1 \\


\ln\left(\frac{y}{z}\right) = C_1 \\

\therefore \frac{y}{z} = C_2, \hspace{1cm} \{\textsf{where}\, C_2 = e^{C_1}\}\\

\textsf{Choosing}\, \left(1, \frac{1}{y\ln(y)}, 0\right)\, \textsf{as multipliers, we have}\\

\mathrm{d}x + \frac{\mathrm{d}y}{y\ln(y)} = 0 \\

\textsf{Integrating both sides, we have} \\
\int \mathrm{d}x + \int \frac{\mathrm{d}y}{y\ln(y)} = C_3\\

x + \int \frac{\mathrm{d}(\ln(y))}{\ln(y)} = = C_3\\


x + \ln(\ln(y)) = C_3\\

\ln(e^x\ln(y)) = C_3\\

e^x \ln(y) = C_4 \hspace{1cm} \{\textsf{where}\, C_4= e^{C_3}\}\\

\textsf{So, the solution of the given PDE is}\, \\ \phi\left(e^x \ln(y),\,\frac{y}{z}\right) = 0"